OpenStudy (anonymous):

Consider a particle with initial velocity that has magnitude 12.0 m/s and is directed 60.0 degrees above the negative x axis. What is the x component Vx of V? What is the y component Vy of V?

OpenStudy (hunus):

 The x component Vx of V is the velocity vector V projected onto the x axis. We obtain this magnitude as follows \[v_{x}=12\cos(120)=-12\cos(60)=-12(\frac{1}{2})=-6\] The y component Vy of V is the velocity vector V projected onto the y axis. We obtain this magnitude as follows \[v_{y}=12\sin(120)=12\sin(60)=12\frac{\sqrt{3}}{2}=6\sqrt{3}=10.4\]